### v2

# Apollo: An Lightweight Multilingual Medical LLM towards Democratizing Medical AI to 6B People

Abstract
Despite the vast repository of global medical knowledge predominantly being in English, local languages are crucial for delivering tailored healthcare services, particularly in areas with limited medical resources. To extend the reach of medical AI advancements to a broader population, we aim to develop medical LLMs across the six most widely spoken languages, encompassing a global population of 6.1 billion. This effort culminates in the creation of the ApolloCorpora multilingual medical dataset and the XMedBench benchmark. In the multilingual medical benchmark, the released Apollo models, at various relatively-small sizes (i.e., 0.5B, 1.8B, 2B, 6B, and 7B), achieve the best performance among models of equivalent size. Especially, Apollo-7B is the state-of-the-art multilingual medical LLMs up to 70B. Additionally, these lite models could be used to improve the multi-lingual medical capabilities of larger models without fine-tuning in a proxy-tuning fashion. We will open-source training corpora, code, model weights and evaluation benchmark.
Figure 1: Linguistic areas covered by Apollo that supports English, Chinese, French, Spanish, Arabic, and Hindi.
1 Introduction
The integration of medical knowledge and artificial intelligence has always been a focal point of research communities, with each incremental improvement potentially enhancing patient experiences and healing rates—serving as a direct manifestation of technology for good. Although medical large language models are promising, existing works are mainly in Chinese (Chen et al., 2023a; Zhang et al., 2023; Bao et al., 2023) or English (Wu et al., 2023; Chen et al., 2023b). The multilingual adaption of medical LLMs, as part of the democratization of large models, seeks to extend the benefits of cutting-edge LLMs to a broader spectrum of users, including those from underrepresented communities. This movement is akin to the historical endeavors to disseminate transformative technologies like electricity and vaccines to wider communities, positing LLMs as the modern equivalents of these essential innovations.
Rationale of Multilinguality in Medical LLMs The rationale of multilingual adaption in medical LLMs (Cox & Maryns, 2021; Pecina et al., 2014) could be twofold. Firstly, non-native English-speaking doctors often engage in bilingual learning through their native language and English from the outset, naturally introducing multilingual challenges in the learning process (Markó et al., 2006). Secondly, to better serve local communities, especially in countries and regions with scarce medical resources, medical aid based on local languages often achieves higher communication efficiency and acceptance (Brindley et al., 2014; Albrecht et al., 2013). Meanwhile, local medical knowledge can complement mainstream medical knowledge, fostering mutual benefits to accelerate medical development (Klayman, 1985; Yuan et al., 2016). Our pilot study in Sec. 2 also reveals that joint training of multiple languages enhances performance in the medical LLMs, indicating a beneficial complementarity among languages.
The Corpora: ApolloCorpora Towards building multilingual medical LLMs, the first step is to build high-qaulity corpora. We select the six most populous languages: English, Chinese, Hindi, Spanish, French, and Arabic for experiments  [^1] See the language popularity in https://en.wikipedia.org/wiki/List_of_languages_by_total_number_of_speakers, The set of the selected six languages covers a total of 6.1 billion people in 132 countries and regions, according to https://en.wikipedia.org/wiki/List_of_countries_and_dependencies_by_population.. As shown in Fig. 2, it encompasses a diverse array of linguistic backgrounds, particularly in regions often characterized by limited medical resources (e.g., areas with lower average life expectancies). We collect and process data from data sources including books, clinical guidelines, encyclopedias, papers, online forums and examinations, obtaining ApolloCorpora with 2.5B tokens.
The Lite LLM: Apollo The resulted multilingual medical LLMs trained by ApolloCorpora are named Apollo, to commemorate the Greek deity associated with healing, disease, the Sun, and light; this symbolizes the democratization of medical LLMs to 6 billion people, illuminating global healthcare. We explore a new domain adaption method that rewrite the pre-training corpora into QA pairs using ChatGPT (Li et al., 2023c) and adaptively sample training data, resulting in a smoother transition compared with the typical paradigm with continued pretraining and instruction tuning. Apollo ranges from 2B to 7B parameters. The advantage of the relatively-small model scale includes potential use as draft models for speculative decoding (Leviathan et al., 2023) or as proxy models for ProxyTuning (Liu et al., 2024a). In particular, we apply proxy tuning on top of Apollo to larger general LLMs, enhancing its multilingual medical capabilities. This is achieved without the need to directly train the general model using sensitive medical corpora, thereby underscoring the practical significance of Apollo in terms of protecting the privacy of medical training data against centralized training methods.
Figure 2: Apollo Pipeline Overview
The Benchmark: XMedBench We select local multiple-choice tasks to assess models’ medical knowledge. For Hindi and Arabic, which lack local assessments, we choose to translate the medical-related parts of MMLU (Hendrycks et al., 2020). The results show that the gap between open source and closed source is narrowing. While GPT-4 has demonstrated superior efficacy across numerous languages, the Apollo series models achieve the best performance among models of equivalent size.
Contributions The contributions of this paper are as follows. (1) We collect and organize a high-quality multilingual medical corpora ApolloCorpora with rich local characteristics; (2) We obtain a series of SOTA multilingual LLMs Apollo at various parameter scales (especially in relatively-small sizes) ; (3) By using proxy tuning, Apollo could significantly improve larger general LLMs without finetuning, providing a new way to mitigating the exposure of private medical training data to centralized training systems; (4) We introduce a multilingual medical evaluation XMedBench and conduct extensive benchmark of existing models.
2 Pilot Study on the Multilinguality of Medical LLMs
2.1 The Research Question
This section presents two contrasting hypotheses regarding the nature of medical knowledge and its representation in LLMs.
Language-neutral Hypothesis It is commonly believed that knowledge, whether medical or general, should be independent of language. For example, the fact that the sun rises in the east remains unchanged whether expressed in English or Chinese, suggesting that knowledge might be considered language-neutral. Consequently, medical corpora in various languages could serve as an augmentation for training, thereby enhancing the efficacy of the resulting medical LLMs.
Language-dependent Hypothesis However, due to historical, cultural, and regional political influences, medical knowledge can vary significantly across different cultural contexts, especially as reflected in language. The integration of medical knowledge across languages might dilute the local specificity of medicine due to differences in lifestyle and constitution across regions (Rotti et al., 2014; Sharma et al., 2020). For instance, in traditional Chinese medicine, colds are classified into types caused by heat or cold, and treatments may vary locally, relying on unique medical recipe that uses local herbs or substances. This indicates that historical and cultural traditions can shape medical knowledge to some extent.
Consequently, there arises a research question in the context of medical LLMs:
Can medical data in different languages complement or harm each?
This leads us to explore whether medical corpora in different languages supplement each other or conflict within medical LLMs.
2.2 The Pilot Study
Model English Chinese French Spanish Arabic Hindi Avg.
Base Model
Qwen-1.8B 32.91 40.07 22.12 27.43 23.71 8.82 25.84
Language Specific Models
Apollo-English 39.44 45.27 28.35 31.76 22.61 8.72 29.36
Apollo-Chinese 39.42 61.13 28.97 33.83 27.94 25.34 36.11
Apollo-French 30.94 32.71 21.81 27.00 24.54 1.74 23.12
Apollo-Spanish 33.84 43.81 27.41 35.39 28.40 23.88 32.12
Apollo-Arabic 36.40 44.27 3.74 15.73 15.90 3.03 19.85
Apollo-Hindi 25.18 3.45 18.38 19.69 1.00 25.53 15.54
Our Method
Apollo (weight average) 40.54 45.58 28.04 34.08 28.95 24.06 33.54
Apollo (multilingual training) 45.43 62.93 38.01 42.15 34.74 25.62 41.48
Table 1: The pilot study on monolingual training and multilingual training.
Experimental settings To investigate the above question, we design a experimental protocol including the following three groups:
  • The backbone LM. We use a lite multilingual LLM Qwen 1.8B (Bai et al., 2023) as the LM backbone. It is argued that the findings should be agnostic to the selection of LM backbones, the selection of LLM Qwen 1.8B is due to its popularity, performance and more the importantly multilingual support.
  • Monolingual training. The LM backbone is further trained by individual language, resulted six language-specific LLM variants (i.e. , English, Chinese, French, Spanish, Arabic, and Hindi.). The training data used could be found in Sec. 3. Moreover, we average the weights of these six LLM variants and obtain a new model that does not need further training, denoted as ‘weight average’.
  • Multilingual training. We train the LM using a mixture of the corpora in these six languages, denoted as ‘multilingual training’.
The evaluation protocol will be detailed in Sec. 4, which shows the average accuracy among datasets in each language.
Findings. Tab. 1 highlights the effectiveness of our methods in leveraging multilingual data to enhance the performance of medical Large Language Models (LLMs). The language-specific models, each trained exclusively on data from one language, demonstrate varying degrees of improvement in their respective languages over the original LM, underscoring the value of language-specific training. However, these models show limitations outside their target languages, as seen in the relatively low scores in non-target languages, particularly in the Apollo-French-1.8B and Apollo-Hindi-1.8B models. Our method, which includes both the weight average and multilingual training, significantly outperforms language-specific models across all languages in terms of average performance (see the last column). This illustrates the substantial benefits of combining multilingual data for training medical LLMs, with marked improvements in understanding and generating medical information across a diverse set of languages.
Therefore, we conclude the finding as below.
In general, multilingual medical corpora benefits medical LLMs.
Potential Risks of Multilingual training in medical LLMs. While acknowledging the potential for conflicts arising from integrating language-specific medical knowledge in multilingual training, we recognize this as a risk inherent in such an approach. However, based on the average performance improvements observed in Tab. 1, we are inclined to believe in the efficacy of multilingual training, especially in the context of medical knowledge, which we argue to be language-neutral to a significant extent. We propose that the conflicts or the potential undermining of local specificities observed in multilingual training be considered as an area for future research. This perspective invites further exploration into how multilingual LLMs can be optimized to respect and preserve the unique medical practices and knowledge embedded within each language, while still harnessing the collective benefits of a multilingual approach.
3 ApolloCorpora, Multilingual Medical Dataset
In this chapter, we first introduce the organizational philosophy of the data set, then introduce the structure of the data set and the data collection process, then introduce the further processing of the data set, and finally analyze the local medical characteristics reflected in the data set.
3.1 Philosophy
Our goal is to enhance the multilingual medical capabilities of large-scale pre-trained models at an affordable computational cost, provide the community with a lightweight reproducible solution, and promote further exploration by the scientific research community. Considering that the model pre-training was most likely trained on PubMed-related data sets, we did not follow the mainstream method and repeat training on all Pubmed data.
If the model that has undergone large-scale pre-training is regarded as a medical student who is about to enter college with certain basic knowledge, in order to obtain more professional capabilities, we only need to replicate the high-quality data learned on the growth path of medical students. Starting from this point, after extensive communication with doctors and medical students, we identified six high-quality medical data collections: medical books, medical encyclopedias, medical clinical guidelines, medical papers, medical examinations, and professional doctor-patient dialogues. At the same time, in order to simulate the data input in the growth process of medical students in addition to medical professional knowledge, we also sampled medical-related parts of the Internet and collected instruction following data for some general tasks (including mathematical tasks of code).
For the dimension of multilingual medical expertise, we insist on using only medical data sets entirely from local languages and do not translate any medical-related data. This is done out of the following two considerations. First, there are many related works that prove that medical translation is a very complex task that cannot be simply solved by translation software; second, the expression habits of different languages, effective drugs, and even culture and Taboo terms arising from faith need to come from the local community intact, so as to maximize communication efficiency and avoid conflicts.
Regarding the License issue, we only screen data sets with complete open source protocols during the collection process to ensure that the open source protocols are friendly while ensuring quality.
Data Source Training Stage Language (# Token) # Token
High-quality medical data
Books Continue pre-training EN (296.7M), ZH (117.1M) 413.8M
Papers Continue pre-training ZH (45.6M), EN (252.9M), ES (46.0M), FR (4.5M) 349.0M
Encyclopedias Continue pre-training EN (221.1M), FR (4.6M), HI (0.5M) 226.2M
Dialogues Continue pre-training EN (92.1M), ZH (46.6M), AR (10.4M) 149.1M
Exams Instruction Tuning EN (42.1M), ZH (35.3M), FR (0.1M), ES (0.5M) 78.0M
Guidelines Continue pre-training EN (29.6M) 29.6M
Data entry outside the profession
Web Continue pre-training EN (499.9M), ZH (329.3M), ES (57.5M) 886.7M
General Instruction Tuning EN (194.5M), ZH (69.4M), HI (43.9M), FR (20.0M), AR (18.7M), ES (18.4M) 364.9M
Math Instruction Tuning EN (18.9M), ZH (3.7M) 22.6M
Code Instruction Tuning EN (9.2M), ZH (7.2M) 16.4M
Table 2: Taxonomy of ApolloCorpora and Token statistics
3.2 Dataset Taxonomy and Collection
As shown in Tab. 2, we collect multilingual data from the data collection direction described in the first section of this chapter, which we will introduce in detail below.
Books
For English books, we use medical dictionary  [^2] https://www.nlm.nih.gov/research/umls/new_users/online_learning/LEX_001.html to filter the books in the Pile Dataset (Gao et al., 2020) and select books with medical words accounting for more than 4%, and finally obtain 2312 medical-related books. For Chinese books, we follow MedQA (Jin et al., 2020) to collect medical textbooks included in the five-year and eight-year medical student training programs in mainland China, and finally obtain 90 books.
Papers
For English papers, we sample the public data in PubMed and obtain 878,241 medical abstracts. For Chinese papers, we also screen a total of 177,261 abstracts of papers published by the Chinese Medical Association  [^3] https://www.yiigle.com/index. For French papers, we use the MORFITT (Labrak et al., 2023b) dataset and the scientific article portion of the CLEAR (Grabar & Cardon, 2018). For the Spanish paper, we use paper abstracts open sourced by the Mesinesp (Gasco et al., 2021).
Encyclopedias
For the English Encyclopedia, we also use the English Medical Dictionary to filter out 36107 medical-related wiki pages from dataset [^4] https://huggingface.co/datasets/wikipedia. For the French encyclopedia, we select the encyclopedia articles part of the CLEAR (Grabar & Cardon, 2018). For the Hindi encyclopedia, we choose the HHD corpus (Jain & Arora, 2018), which crawls descriptions of people, diseases, medical consumer products, and symptoms from Indian websites.
Doctor-Patient Dialogues
For Chinese, we directly use the HuatuoGPT dataset (Zhang et al., 2023) and the simplified data set in Huatuo_26M (Li et al., 2023b). For English, we construct a multi-turn conversation data set based on PMC-Patients (Zhao et al., 2022) using ChatGPT, Prompt is shown in the Appendix A.1. For Arabic, we extract high-quality questions and answers with both question and answer lengths greater than 128 from the largest Arabic healthcare question and answer dataset MAQA (Abdelhay & Mohammed, 2022).
Exams
For the Chinese exam, we collect training sets of CMB (Wang et al., 2023), CMExam (Liu et al., 2024b), and MedQA (Zhang et al., 2018). For the English exam, we collect the training sets of MedQA, Medmcqa (Pal et al., 2022) and Pubmedqa (Jin et al., 2019). For the Spanish and French exam, we select the training set of HEAD-QA (Vilares & Gómez-Rodríguez, 2019) and Frenchmcqa (Labrak et al., 2023a) separately.
Guidelines
For English Guidelines, we select data from three sub-items of NICE [^5] https://www.nice.org.uk/guidance, PubMed and SPOR [^6] https://sporevidencealliance.ca/key-activities/cpg-asset-map/cpg-database/ in the clinical guidelines introduced by Meditron (Chen et al., 2023b).
General Instruction Tuning
We use the translation (FreedomIntelligence, 2023a) and original data of Sharegpt [^7] https://github.com/lm-sys/FastChat and Alpaca (Taori et al., 2023). For Chinese, we additionally make use of data (FreedomIntelligence, 2023b) generated by GPT-4 based on WizardLM Method (Xu et al., 2023). For English, in addition to adding the WizardLM Dataset, we also add belebele (Bandarkar et al., 2023) to enhance multi-language reading comprehension capabilities, ai2_arc (Clark et al., 2018) to enhance abstract reasoning capabilities, Capybara (Daniele & Suphavadeeprasit, 2023) to enhance instruction following capabilities.
Web
For Chinese, we use the medical dictionary (Han et al., 2016) to filter out medical-related articles from the Wudao Dataset (Xue et al., 2022). For English, we use the English Medical Vocabulary [^8] https://www.nlm.nih.gov/research/umls/new_users/online_learning/LEX_001.html to filter out medical related articles in C4 Dataset (Raffel et al., 2019). For Spanish, we sampled 10% of CoWeSe Dataset (Carrino et al., 2021).
Math
For mathematical abilities, we choose MathInstruct (Yue et al., 2023), a composite dataset containing various mathematics-related tasks and problem formats.
Code
We choose Python-Alpaca (Vezora, 2023) and Leetcode-ZH-11k (krisfu, 2023) respectively to strengthen the ability to solve coding tasks in Chinese and English.
3.3 Data processing
Rewrite
If we continue to train plain text, it is likely to destroy the original output format and knowledge. We want to explore whether rewriting the original pre-training corpus into QA pairs in the context of continuing training can help increase its medical capabilities without destroying the original model’s capabilities. We use ChatGPT999gpt-3.5-turbo-16k-0613 to generate questions and answers for a certain paragraph. For paragraph interception, we divide it according to the basic semantic units in the data set, such as sections in books and guides, paragraphs in website data, single wiki entry and abstracts of papers. For basic semantic units that are too long, we comprehensively consider the knowledge expression density of the language and subdivide different languages into blocks of different lengths to ensure that the semantic information covered by a single paragraph does not exceed the amount of information that can be included in a question and answer pair. For Spanish, French, English and Hindi we use 2048, for Chinese we use 256 and for Arabic we use 128. Prompts for generating QA pairs are detailed in the Appendix A.2.
Data Leakage Checking
The issue of data leakage is a recent focus of the academic community, which largely determines whether the results of the paper are convincing. For knowledge embedding tasks, data leakage screening with different stringency often leads to different performance. We follow Med-PaLM2 (Singhal et al., 2023) and adapt a more stringent deletion strategy. Specifically, we define a data item as leaked data if the entire question or at least 64 consecutive characters overlap with the data item. Regarding the exam exercise data source, there were 580,645 exercises before screening, and 3,041 exercises were deleted, with a screening rate of 0.52%. For other data sources, since they are not exam questions, there is no difference before and after filtering.
Examples Chinese: Representatives of exogenous wind-cold and high fever include: (A) Yinqiao Powder, (B) Qingwen Baidu Decoction, (C) Dachengqi Decoctio. (D) Xiaobuhu Decoction Arabic: Patient: “Peace be upon you, I feel very intense friction in my body and head, redness in the skin or its color changing to pink, and I am almost cut off in my body. This all happens when I start making an effort, even if it is the smallest of things, even if my voice rises or I get angry. Thank you, and may God’s peace, mercy, and blessings be upon you.” Doctor: “Often this is sensitivity due to exertion and sweating, and they are signs of a disease called fever of the nile, or a rash or allergy to sweating that appears in the summer season in particular. It is preferable to use a cold shower and apply the pimples with a weakening ointment. If the area is small, apply Calcipotriene ointment to it, take an anti-histamine medication such as Polaramine.” Hindi: Disease: Epilepsy. Tulsi reduces many diseases like a panacea. A large amount of anti-oxidants are found in Tulsi which cures free radicals in the brain. In case of any type of brain disease, if taken daily If 20 basil leaves are chewed and eaten, it is very effective. Brahmi leaves: Brahmi leaves grow around our homes especially where there is soil. It is round and curved in shape. It is taken daily. Consuming it chewed on an empty stomach not only strengthens the memory but also reduces epileptic seizures. Spanish: Due to its high mercury content, the Spanish Agency for Food Safety and Nutrition recommends pregnant women not consume: A. Ripe soft cheeses, such as Brie or Camembert, B. Pâté or foie-gras, C. Calmette and Guérin bile vaccine., D. Raw sausages French: What vaccinations are required in France? A.Measles-Mumps-Rubella, B.Human papillomavirus, C.Diphtheria-Tetanus-Poliomyelitis, D.Whooping cough, E.Calmette and Guérin bile vaccine.
Figure 3: Examples of local language characteristics in ApolloCorpora
3.4 Localized features
As shown in the Fig. 3, we illustrate the local language features in the dataset by language:
In terms of symptom diagnosis, local languages retain the terminology of traditional medicine, and due to different geographical environments and living habits, the possibility that a certain symptom corresponds to different diseases is also different: for Chinese, a disease has two aspects: ”bìng” and ”zhèng”. The former is often translated as ”disease entity”. The latter, and more important one, is usually translated as ”pattern”. For example, the disease entity of a common cold might present with a pattern of wind-cold in one person, and with the pattern of wind-heat in another [^10] https://en.wikipedia.org/wiki/Traditional_Chinese_medicine#Six_Excesses.
In terms of medicines, each language has its own specific names for medicines, and even retains some medicines from traditional medicine: for Chinese, there are about 13,000 medicines recorded in ancient Chinese literature and more than 100,000 Chinese medicine prescriptions; for Arabic and Hindi, doctors may also include some local plants in their medicines.
In terms of communication terms, some languages will have some religious-related idioms at the beginning and end to improve the communication experience, such as Arabic.
In terms of medical practice standards and dietary recommendations, different medical systems have different standards, and different places also have different customary diets: for Spanish and French, local standards may differ, and dietary recommendations are also consistent with the preferences of the local population.
4 XMedBench, Multilingual Medical Knowledge Assessment
Construction
We focus on assessing multilingual medical knowledge, select multiple-choice questions as assessment tasks, and collect common data sets with local medical characteristics as much as possible. For English, we use the MedQA-USMLE (Zhang et al., 2018), MedMCQA (Pal et al., 2022), and medical-related parts of MMLU (Hendrycks et al., 2020); for Chinese, we used the of MedQA-MCMLE (Zhang et al., 2018) and medical-related parts CMMLU (Li et al., 2023a); for Spanish, we used HEAD-QA (Vilares & Gómez-Rodríguez, 2019); for French, we used FrenMedMCQA (Labrak et al., 2023a); For Arabic and Hindi, which lack local assessments, we make a compromise by applying translated versions of MMLU [^11] Hindi: https://huggingface.co/datasets/FreedomIntelligence/MMLU_Hindi; Arabic: https://huggingface.co/datasets/FreedomIntelligence/MMLU_Arabic. Specifically, we follow Med-PaLM2 (Singhal et al., 2023) and select six subcategories in MMLU: Clinical knowledge, Medical genetics, Anatomy, Professional medicine, College biology, and College medicine. For MedQA, we choose the 4-options version. For CMMLU, we select seven subdirectories: Anatomy, Clinical knowledge, College medicine, Genetics, Nutrition, Traditional chinese medicine, and Virology.
Prompt User:You are a medical doctor answering real-world medical exam questions. Select one correct answer from A to D. Question: {question} Options: (A) {optiona} (B) {optionb} (C) {optionc} (D) {optiond} Assistant:The correct answer is {answer}. <special_token >
Figure 4: Prompt Template for XMedBench
Settings
We adapt 3-shot evaluation, and the specific evaluation prompt is shown in the Fig. 4. To determine the answer, we use regular matching from the answers generated by the model. For the generation strategy, we do not perform sampling and set the maximum and minimum number of generated tokens to 128 and 2. For model loading, except for the 0.5B size model which uses full precision loading, we uniformly use half precision loading. Please see the Appendix for details of the models.
Language English Chinese French Spanish Arabic Hindi Avg.
USMLE MedMCQA MMLUnormal-$\diamondsuit$ MCMLE CMMLUnormal-$\diamondsuit$ FrenchMedMCQA HEAD-QA MMLUnormal-$\diamondsuit$ MMLUnormal-$\diamondsuit$
Closed-source
GPT-4 79.10 70.40 86.00 65.72 65.72 89.72 85.05 56.43 62.17 73.37
GPT-3.5 61.98 56.51 72.94 58.73 50.41 68.54 71.48 39.70 39.94 57.80
Open-source (Above 70B)
Qwen-72B 64.10 62.16 78.46 91.68 81.47 74.14 76.62 46.87 43.16 68.74
Meditron-70B 55.70 50.87 69.59 48.34 40.29 53.27 59.74 19.30 31.31 47.60
Llama-2-70B 32.99 48.29 64.62 25.80 25.13 50.47 54.34 1.65 26.35 36.63
Open-source (Above 7B)
Qwen-14B 50.27 45.83 61.68 75.22 61.82 49.53 60.81 36.58 32.29 52.67
Gemma-7B 53.42 50.94 70.15 48.95 43.29 57.63 62.79 36.21 48.58 52.44
MMedLM2-7B 55.46 50.49 68.15 64.30 56.11 58.57 62.14 23.53 24.15 51.45
Yi-34B 62.45 60.60 71.86 26.12 26.51 66.04 69.99 30.70 9.73 47.00
Mistral-7B 47.29 47.38 62.80 38.32 34.21 50.78 51.93 28.40 27.36 43.16
Qwen-7B 32.36 39.52 53.22 54.32 44.71 37.69 45.05 28.31 24.89 40.01
Zephyr-7B-$\beta$ 41.95 42.48 58.74 36.11 31.88 46.42 46.77 27.02 27.92 39.92
BioMistral-7B 41.79 42.05 54.46 34.65 31.43 43.61 44.66 27.11 22.96 38.08
Huatuo2-7B 37.86 36.58 42.49 55.08 43.81 27.41 33.88 25.92 27.46 36.72
Huatuo2-13B 29.77 36.58 42.86 56.07 45.46 22.42 36.13 18.29 13.59 33.46
Llama-2-7B 32.13 36.58 40.14 25.39 25.13 29.60 33.54 21.42 27.27 30.13
Meditron-7B 33.78 34.54 36.18 27.50 27.16 24.00 32.81 1.65 18.27 26.21
PMC-Llama-7B 20.11 23.12 19.72 16.90 16.73 17.13 18.68 9.65 2.85 16.10
Our Models
Apollo-0.5B 32.99 37.82 45.87 56.57 42.08 27.41 36.67 31.89 25.90 37.47
Apollo-1.8B 42.18 44.99 49.12 72.30 53.56 38.01 42.15 34.74 25.62 44.74
Apollo-2B 38.33 42.00 52.89 46.76 36.76 38.32 41.28 31.62 31.50 39.94
Apollo-6B 56.25 57.53 68.65 85.52 72.62 51.71 58.47 33.46 33.61 57.54
Apollo-7B 56.00 58.21 71.86 72.36 59.04 60.44 63.73 41.82 45.55 58.78
Table 3: Performance comparison across various medical question answering models.
Results
As shown in Tab. 3, GPT-4 and Qwen-72B rank first in closed source and open source with accuracy rates of 73.37 and 68.74 respectively. The gap between closed source and open source is gradually decreasing. The Apollo series models achieve the best performance of models of the same size. Apollo-7B achieve the same performance as GPT-3.5, Apollo-1.8B achieve the same performance as Mistral-7B, and Apollo-0.5B achieve the same performance as Llama2-7B.
From a language perspective, the models generally does not support Arabic and Hindi well, which further reflects the importance of multilingual medical care. It is worth noting that the two models GPT-4 and Gemma-7B clearly support these two languages better, reflecting the optimization of top open source models for multi-language scenarios. For French, Mistral is the best model of its size. For Spanish, performance is roughly in sync with English. For Chinese, Qwen series models perform better.
At the same time, we also discover some interesting phenomena. BioMistral’s continued training on PubMed-Central based on Mistral does not improve significantly, and the same is true for Meditron. This further confirm that continuing to train extensively on PubMed-Central has little effect and low cost performance, which is mentioned in Sec 3.1.
5 Apollo, the Lite Multilingual Medical LLM
We have two main starting points for training the small model. First, medical equipment usually cannot call network services due to its strict privacy protection settings. For local services, the small model can achieve offline inference on the PC side, which is completely Ensure that the data is localized to help improve the efficiency of medical staff; secondly, the original intention of our article is to explore a reproducible technical solution at an affordable computing cost, and promote the exploration of the field and the raising of new questions. Small models are useful for The training is very friendly for academic researchers who lack sufficient computing power.
Training models in the medical field usually involves continuing pre-training on the corpus. However, some scholars believe that although training on the original corpus gives the model domain knowledge, it greatly damages its ability to prompt question answers (Cheng et al., 2023). We consider exploring ways to rewrite the pre-training corpus into the form of question-and-answer pairs to alleviate this problem (Chen et al., 2023a). At the same time, we use priority sampling methods (Chen et al., 2023a) to achieve a smooth transition between continued pre-training and Instrcution Tuning to ensure the continuity of learning rate and data distribution transformation.
Preliminaries
Our dataset $D$ comprises continuing pre-training data $D_{PT}$ and instruction tuning data $D_{SFT}$. The sampling probability of each data $x\in D$ changes over the course of training. The sampling probability of data $x$ at step $t$ during training was determined using priority sampling, defined as:
$P_{t}(x)=\frac{\pi(x)}{\sum_{y\in D-S_{t}}\pi(y)}$
Here, $\pi(x)$ denotes the priority of element $x$, and $S_{t}$ represents the sampled data before step $t$.
Settings
We set the priority of all data items from the pre-training corpus to 16, and the priority of all data items from the Instruction Tuning stage to 2. The Batch size of model training is set to 256, the learning rate is set to 1e-4, and the warm up rate of Cosine scheduler is set to 0.03. The pre-training corpus is trained for one epoch, the Instrument Tuning corpus is trained for two epochs.
Model English Chinese French Spanish Arabic Hindi Avg.
Base Model
Qwen-1.8B 32.91 40.07 22.12 27.43 23.71 8.82 25.84
Rewrite Pre-training Data into QA
ParaData-Sep-1.8B 47.34 57.58 37.69 41.24 29.32 15.79 38.16
QAData-Sep-1.8B 45.43 59.21 38.01 42.48 31.43 14.60 38.53
Smoothly Transition the Two Stages
ParaData-Mix-1.8B 42.97 53.56 33.02 36.88 31.71 14.23 35.40
QAData-Mix-1.8B (Apollo-1.8B) 45.43 62.93 38.01 42.15 34.74 25.62 41.48
Table 4: Mix Training for Multilingual.
Language English Chinese French Spanish Arabic Hindi Avg.
USMLE MedMCQA MMLUnormal-$\diamondsuit$ MCMLE CMMLUnormal-$\diamondsuit$ FrenchMedMCQA HEAD-QA MMLUnormal-$\diamondsuit$ MMLUnormal-$\diamondsuit$
Our Models and their Bases
Qwen-0.5B 24.43 3.78 16.94 14.16 10.88 23.68 26.02 26.29 26.35 19.17
Apollo-0.5B 32.99 37.82 45.87 56.57 42.08 27.41 36.67 31.89 25.90 37.47
Qwen-1.8B 26.79 31.05 40.89 44.28 35.86 22.12 27.43 23.71 8.82 28.99
Apollo-1.8B 42.18 44.99 49.12 72.30 53.56 38.01 42.15 34.74 25.62 44.74
Gemma-2B 30.24 32.27 37.35 25.98 28.06 25.86 32.43 20.96 25.53 28.74
Apollo-2B 38.33 42.00 52.89 46.76 36.76 38.32 41.28 31.62 31.50 39.94
Yi-6B 45.48 47.98 62.27 78.90 69.47 45.79 47.01 12.22 10.74 46.65
Apollo-6B 56.25 57.53 68.65 85.52 72.62 51.71 58.47 33.46 33.61 57.54
Gemma-7B 53.42 50.94 70.15 48.95 43.29 57.63 62.79 36.21 48.58 52.44
Apollo-7B 56.00 58.21 71.86 72.36 59.04 60.44 63.73 41.82 45.55 58.78
Table 5: Model performance comparison before and after Mix Training
Results
As shown in the Tab. 4, under the experimental setting of pre-training first and then SFT, rewriting the pre-training into question and answer has no loss on the overall effect of the model. We also notice that the performance drop after rewriting in English and Hindi, but other languages’ performance improve. After adopting the smooth transition method, we find that except English, other languages’ performance greatly improve. This may be because the data distribution transformation of the previous method is too rigid, resulting in the inability to learn knowledge of long-tail languages (such as Hindi). Using the method of converting to question and answer pairs and making a smooth transition may be able to minimize the knowledge loss of distribution transformation, allowing the model to fully learn the knowledge of long-tail languages and improve the ability of non-mainstream languages. As shown in the Tab. 5, models’ multilingual medical capabilities have been significantly improved after Mix Training.
Model English Chinese French Spanish Arabic Hindi Avg.
Other Models
GPT-3.5 63.81 54.57 68.54 71.48 39.70 39.94 56.34
Meditron-7B 34.83 27.33 24.00 32.81 1.65 18.27 23.15
ProxyTuning for Qwen
Apollo-1.8B (from Qwen-1.8B) 45.43 72.30 38.01 42.15 34.74 25.62 43.04
Qwen-7B 41.70 49.52 37.69 45.05 28.31 24.89 37.86
Qwen-7B-ProxyTuning 39.83 -${}_{\textbf{-1.87}}$ 51.40 ${}_{\textbf{+1.88}}$ 43.30 ${}_{\textbf{+5.61}}$ 46.97 ${}_{\textbf{+1.52}}$ 29.69 ${}_{\textbf{+1.38}}$ 24.89 ${}_{\textbf{+0.00}}$ 40.79 ${}_{\textbf{+2.93}}$
Table 6: Proxy Tuning for Larger Models
6 The application of the lite Apollo: Proxy Tuning for Larger Models
Preliminaries
Inspired by  Liu et al. (2024a; 2021), we introduce a lightweight model-agnostic decoding method in medical senarios. We leverage the logits from both pre and post fine-tuned small models to indirectly steer the larger base model’s adjustments, thereby eschewing the need for direct parameter fine-tuning. Let $M_{raw}$ denote the smaller pre-trained model, and $M_{tuned}$ denote its fine-tuned counterpart. We compute the logit offset as ”proxy” for each token, corresponding to the anti-expert and expert roles as delineated in  Liu et al. (2021). This offset is then applied to the base model $M_{base}$ to synchronize the predictive distributions of the smaller and larger models. The modified probability distribution is given by:
$p^{\prime}_{\mathcal{M}_{base}}(X_{t}\mid x_{1,...,t-1})=\text{softmax}\left[l_{M_{base}}+\Delta l_{M}\right]\propto p_{\mathcal{M}_{base}}(X_{t}\mid x_{1,...,t-1})\left(\frac{p_{\mathcal{M}_{tuned}}(X_{t}\mid x_{1,...,t-1})}{p_{\mathcal{M}_{raw}}(X_{t}\mid x_{1,...,t-1})}\right)$
where $\Delta l_{M}=l_{M_{tuned}}-l_{M_{raw}}$ represents the logit offset of the expert model $M_{tuned}$ and the anti-expert pre-trained model $M_{raw}$.The logit output for $M$ at $t$ is denoted by $l_{M_{t}}$ for the current timestep $t$. The probability distribution of $M$ refers to $p_{\mathcal{M}_{base}}(X_{t}\mid x_{1,...,t})$.
Settings
$M_{base}$ is designated as the subject of investigation for Qwen-7B. Apollo-1.8B and Qwen-1.8B are appointed as the $M_{tuned}$ and $M_{raw}$ respectively.
Results
As shown in the Tab. 6, the overall effect of the model improves a lot without changing the parameters after Proxy Tuning. From language perspective, except English, all other languages increase, and French has the most obvious increase. Excitingly, for French and Spanish, the model after Proxy Tuning performs better than both $M_{tuned}$ and $M_{base}$, indicating that new accurate knowledge is generated after Proxy Tuning. We also notice a decline in English proficiency. This may be because there is a gap between the distribution of difference and the probability itself, which leads to over-strengthening of the second option and requires further exploration and optimization.
7 Related Work
[…]
Notable early developments in healthcare-focused LLMs include GatorTron (Yang et al., 2022), which explores the use of unstructured Electronic Health Records (EHRs) in clinical settings, and Codex-Med (Liévin et al., 2022), which assesses the efficacy of GPT-3.5 models in healthcare. Galactica (Taylor et al., 2022) targets the challenge of scientific information overload by storing, combining, and reasoning about scientific knowledge, including in healthcare.
Privacy and confidentiality concerns have prompted innovative solutions like DeID-GPT (Liu et al., 2023), a de-identification framework using GPT-4, and MedAlpaca (Han et al., 2023), which offers an open-source, on-site implementation approach. ChatDoctor (Yunxiang et al., 2023), Meditron (Chen et al., 2023b) and PMC-LLaMA (Wu et al., 2023) address limitations in medical knowledge accuracy of existing LLMs, tuning base models on millions of biomedical papers.
GatorTronGPT (Peng et al., 2023), a clinical LLM with a GPT-3 architecture, and MedAGI (Zhou et al., 2023), a model aiming to unify domain-specific medical LLMs, illustrate the trend towards more specialized models. Finally, Med-Flamingo (Moor et al., 2023) represents an advance in vision-language models, handling multimodal data that includes both images and text.
There have been some outstanding works focusing on multilingual topics recently. BioMistral (Labrak et al., 2024) introduce the perspective of a multilingual evaluation system for the first time. MMedLM (Qiu et al., 2024) is the first large medical model trained on multilingual corpus. We believe that our work, together with the formers, will bring a multilingual perspective into the medical artificial intelligence community and help more people with Medical AI.
8 Conclusion
In order to serve more people and larger community, we carefully collect and organize a high-quality medical corpus covering most populous languages in the world, open sourcing multi-language Dataset ApolloCorpora and evaluation set XMedBench. Based on these, we explore suitable methods for multilingual training and interrelationships between languages in the medical field, and finally obtains a series of models named after Apollo, with SOTA performance from 0.5B to 7B. At the same time, the Proxy Tuning method is also used to improve larger models’ multilingual medical capabilities without changing the parameters.
We propose the question of exploring the relationship between languages in the medical field. While open-sourcing a series of small models with comparable performance, we expand their application scenarios and keep the entire process open-source and reproducible. We provide a foundation for researchers with limited resources to explore questions such as datasets sampling optimization, Proxy Tuning optimization, and different language models combination optimization in the multilingual medical field.
Acknowledgements
This work is supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), and Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608).
We recognize the existence of a concurrent work by Qiu et al. (2024) (Qiu et al., 2024), which became publicly available prior to our work. It is important to note, however, that our research was conducted before the release of Qiu et al.’s work, with the interval before our publication being devoted to thorough review and refinement of our manuscript.
Appendix A Prompts
A.1 Prompt for Generating Doctor-Patient Dialogues
Prompt <text>{text}</text> Please create some dialogues between patients and doctors in English based on the above text. The format is: <Patient>Patient’s question</Patient> <Doctor>Doctor’s answer</Doctor> Both patient questions and doctor responses are as complex and detailed as possible.
Figure 5: Prompt Template for Generating Doctor-Patient Dialogues
A.2 Prompts for Generating QA Pairs from Texts
Prompts Prompt for Generating Question: Please create a <question >that closely aligns with the provided <text>. Ensure that the <question>is formulated in English and does not explicitly reference the text. You may incorporate specific scenarios or contexts in the <question>, allowing the <text>to serve as a comprehensive and precise answer. <text>: {text} <question>: Prompt for Generating Answer: You are Apollo, equipped with in-depth knowledge in medicine. Your task is to directly answer the user’s <question>in English. In formulating your response, you must thoughtfully reference the <reference text>, ensuring that your reply does not disclose your reliance on <reference text>. Aim to provide a comprehensive and informative response, incorporating relevant insights from <reference text>to best assist the user. Please be cautious to avoid including any content that might raise ethical concerns. <question>: {question} <reference text>: {reference} <reply>:
Figure 6: Prompts for Generating QA Pairs from Texts. We show the English version of Prompt, and other languages are similar.
Appendix B Models
Qwen
Qwen is a suite of large language models from the Aliyun-developed Tongyi Qianwen from 0.5 billion to 72 billion parameters, based on the Transformer architecture and are trained on a diverse and extensive range of pretraining data. The types of pretraining data are varied and cover a wide scope, including a vast array of internet texts, professional books, code, and more.
Meditron
Meditron is a suite of open-source medical large language models from 7 billion to 70 billion parameters, adapted to the medical domain from Llama-2 through continued pretraining on a comprehensively curated medical corpus, including selected PubMed articles, abstracts, a new dataset of internationally-recognized medical guidelines, and general domain data from RedPajama-v1.
Llama-2
Llama-2 is a collection of pretrained and fine-tuned generative text models ranging in scale from 7 billion to 70 billion parameters. The fine-tuned versions use supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) to align to human preferences for helpfulness and safety.
Gemma
Gemma is a family of lightweight, state-of-the-art open models from Google, built from the same research and technology used to create the Gemini models. They are text-to-text, decoder-only large language models, available in English, with open weights, pre-trained variants, and instruction-tuned variants.
MMedLM2
MMedLM 2 is a multilingual medical foundation model available in two versions, with parameter sizes of 1.8 billion and 7 billion. MMedLM 2 builds upon the foundation of InternLM 2 and has been further pretrained on MMedC, a comprehensive multilingual medical corpus. This further pretraining enhances the model’s medical-domain knowledge.
Yi
The Yi series models are the next generation of open-source large language models trained from scratch by 01.AI. Targeted as a bilingual language model and trained on 3T multilingual corpus, the Yi series models show promise in language understanding, commonsense reasoning, reading comprehension, and more.
Mistral
Mistral is a pretrained generative text model with 7 billion parameters. It uses Grouped-query attention (GQA) for faster inference and Sliding Window Attention (SWA) to handle longer sequences at smaller cost.
Zephyr
Zephyr is a series of language models that are trained to act as helpful assistants, which is a fine-tuned version of mistralai/Mistral-7B-v0.1 that was trained on on a mix of publicly available, synthetic datasets using Direct Preference Optimization (DPO).
BioMistral
BioMistral is a suite of Mistral-based further pre-trained open source models suited for the medical domains and pre-trained using textual data from PubMed Central Open Access. All the models are trained using the CNRS (French National Centre for Scientific Research) Jean Zay French HPC.
HuatuoGPT-2
HuatuoGPT2 is a suite of open-source medical large language models from 7 billion to 34 billion parameters, which employs an innovative domain adaptation method to significantly boost its medical knowledge and dialogue proficiency. It showcases state-of-the-art performance in several medical benchmarks, especially surpassing GPT-4 in expert evaluations and the fresh medical licensing exams.
PMC-Llama
MedLlama is initialized from Llama-13B and further pretrained with medical corpus. Despite the expert knowledge gained, it lacks instruction-following ability. It provides a instruction-tuning dataset and evaluates the tuned model. MedLlama is pretrained on medical corpus, and PMC_Llama is further finetuned based on MedLlama.
Apollo (Ours)
Apollo is a suite of open-source medical large language models from 1.8 billion to 7 billion parameters. The priority of all data items from the pre-training corpus to 16, and the priority of all data items from the Instruction Tuning stage to 2. The Batch size of model training is set to 256, the learning rate is set to 1e-4 for most models and 1e-5 for 7B model, and the warm up rate of Cosine scheduler is set to 0.03. The pre-training corpus is trained for one epoch, the Instrument Tuning corpus is trained for two epochs.